The student decides to determine the molarity of the same Na2CO3 solution using a second method. When Na2CO3 is dissolved in water, CO3 ^2−(aq) hydrolyzes to form HCO3 ^−(aq), as shown by the following equation.

CO3 2−(aq) + H2O(l)  HCO3 −(aq) + OH−(aq)    Kb = [HCO3^ -][OH^- ]/ [CO3^2- ] - - - = 2.1 × 10^−4

(f) The student decides to first determine [OH−] in the solution, then use that result to calculate the initial concentration of CO3 2−(aq).

(i) Identify a laboratory method (not titration) that the student could use to collect data to determine [OH−] in the solution.

(ii) Explain how the student could use the measured value in part (f)(i) to calculate the initial concentration of CO3 2−(aq). (Do not do any numerical calculations.)

(g) In the original Na2CO3 solution at equilibrium, is the concentration of HCO3 −(aq) greater than, less than, or equal to the concentration of CO3 2−(aq) ? Justify your answer.

(h) The student needs to make a CO3 2−/HCO3 − buffer. Is the Na2CO3 solution suitable for making a buffer with a pH of 6? Explain why or why not.
Answer #1

The answer is attach below

ANSWER! CORE) + H2O G + HCG @u tohºijer) kn = [HC03][0m [co] Kb = 2.1 x64, is À laboratory method (note Hitration that the stHence (20,43 = Cowg from pl molbiz Note: (Coz? I = m from given value. concentraction at equilibriun we know [Coz?] = @ initi41 Kb = 2.1x10-4 Kaxk6=1514 ka=1614 = 4.76 x 10 2- 2x 10-4 = pka=-1 og [ka] pka : - log (4.76X4011) = 10.3222 No, the student